OpenStudy (some.random.cool.kid):

Factor 7(x − 3)2 − 4(x − 3) − 3 completely.

OpenStudy (some.random.cool.kid):

any ideas?

OpenStudy (some.random.cool.kid):

@mathstudent55 @zepdrix @uri

OpenStudy (some.random.cool.kid):

@sleepyjess

OpenStudy (some.random.cool.kid):

@StudyGurl14

OpenStudy (some.random.cool.kid):

@Jaynator495 @just_one_last_goodbye

OpenStudy (some.random.cool.kid):

@tanya123 @ikram002p

zepdrix (zepdrix):

Hey :)

OpenStudy (some.random.cool.kid):

zep do you know this?

zepdrix (zepdrix):

\[\large\rm 7(x − 3)^2 − 4(x − 3) − 3\]Is the -3 supposed to be an exponent as well?

OpenStudy (some.random.cool.kid):

i have a few of these to do... but im just starting with this one.

OpenStudy (some.random.cool.kid):

mmm no i dont believe so

zepdrix (zepdrix):

Let's write it like this\[\large\rm =7(x-3)(x − 3) − 4(x − 3) − 3\]

zepdrix (zepdrix):

Notice that the first and second terms have something in common,\[\large\rm =7(x-3)\color{orangered}{(x − 3)} − 4\color{orangered}{(x − 3)} − 3\]

OpenStudy (some.random.cool.kid):

yes

OpenStudy (some.random.cool.kid):

so do i cancel them?

OpenStudy (some.random.cool.kid):

=7(x - 3) - 4 - 3? if canceling like terms correct.

zepdrix (zepdrix):

Maybe we should put some brackets here, just to prepare ourselves for what we'll do next,\[\large\rm =\left[7(x-3)\color{orangered}{(x − 3)} − 4\color{orangered}{(x − 3)}\right] − 3\] So we divide out this common factor, and leave it as multiplication on the outside like this.\[\large\rm =\color{orangered}{(x − 3)}\left[7(x-3) − 4\right] − 3\] Yes, you have the right idea :) Divide (x-3) out of each term, but you can't just divide willy nilly. To keep things balanced, you have to also multiply at the same time. So we're dividing an (x-3) out of both terms, and saving the multiplication on the outside. If that's too confusing, we can do a more simple example.

OpenStudy (some.random.cool.kid):

no, no I see your point. However I am confused as to which i multiply?

OpenStudy (some.random.cool.kid):

@zepdrix

zepdrix (zepdrix):

You only multiply (x-3) by the terms which you took an (x-3) from. So notice that we didn't include the -3 in the square brackets. We didn't "factor" an (x-3) out of that, so we don't want him to be included in the multiplication. \[\large\rm =(x − 3)\left[7(x-3) − 4− 3\right]\qquad\qquad \color{red}{X}\]\[\large\rm =(x − 3)\left[7(x-3) − 4\right] − 3\qquad\qquad \color{green}{\checkmark}\]

zepdrix (zepdrix):

Is that what your question was about? :o Maybe I misunderstood.

OpenStudy (some.random.cool.kid):

ok so what shall i d next?

zepdrix (zepdrix):

Hmm, I guess you could simplify the stuff that's inside of the square brackets. \[\large\rm =(x − 3)\left[\color{royalblue}{7(x-3) − 4}\right] − 3\]

OpenStudy (some.random.cool.kid):

ok so to simplify what must be done? just multiply divide? or do i factor?

OpenStudy (some.random.cool.kid):

oh i see your point? multiply the 7 by whats in the brackets?

zepdrix (zepdrix):

Ya distribute the 7 :) that's a good idea

OpenStudy (some.random.cool.kid):

hmm so I see 7x - 21) - 28?

zepdrix (zepdrix):

Woops! Notice that the 7 is only multiplying the (x-3), not the -4.

OpenStudy (some.random.cool.kid):

so 7x - 21? or is it 21x?

zepdrix (zepdrix):

You give a 7 to the x, and a 7 to the -3. So 7x and -21, yes. \[\large\rm =(x − 3)\left[\color{royalblue}{7x-21 − 4}\right] − 3\]

OpenStudy (some.random.cool.kid):

sorry internet issues...

OpenStudy (some.random.cool.kid):

so now what? 21 - 4? then (7x - 17) is left right?

zepdrix (zepdrix):

Woops! They're both negative, so you end up with a number which is more negative.

zepdrix (zepdrix):

-21 and go 4 more negative.

OpenStudy (some.random.cool.kid):

uh so -25?

zepdrix (zepdrix):

\[\large\rm =(x − 3)\left[\color{royalblue}{7x-25}\right] − 3\]K cool. We can change those back to round brackets at this point if you want.\[\large\rm =(x − 3)\left(7x-25\right) − 3\] There is no difference in bracket, it was just so we wouldn't have a (( situation :) Easier to read, I think. I guess we can't do much to it besides all of that. :o

OpenStudy (some.random.cool.kid):

so more distributing?

zepdrix (zepdrix):

No, nothing more. That's all the factoring we can do :) Distributing would be going the other direction.

OpenStudy (some.random.cool.kid):

uh ok

OpenStudy (some.random.cool.kid):

can i make a new tab this is starting to have connection issues.

OpenStudy (some.random.cool.kid):

@zepdrix i will make new tab and update you.

zepdrix (zepdrix):

k :)